Can someone hlep me solve this problem?
I need to multiply:
(-3) in the power of 3 - (-8)(-2)
Thanks,
Lyn

Answer by fractalier(2101)

This time we have
(-3)^3 - (-8)(-2) =
(-27) - (16) =
-27 - 16 =
-43